# Using Biot-Savart boundary conditions for unbounded external flow on Eulerian meshes

### v2

@@ 5.2 Wake behind a heaving foil
We immerse a rigid foil of length $L$ moving with a linear velocity $\vec{U}=(U,0)$ in a viscous fluid. In addition to its steady forward speed, the foil undergoes a pure heave motion of amplitude $h_{0}$
@@
where $h_{0}/L=0.5$ is the non-dimensional amplitude, and the frequency is $f=\text{St}\,U/h_{0}$. For Strouhal number $\text{St}<0.5$, the airfoil generates thrust, and almost no net lift, and for $\text{St}>0.6$, the wake generated by this flapping foil is strongly asymmetric and generates a significant net lift force. We perform simulations of this system for a range of Strouhal numbers and compare the mean lift force and wake obtained on a large domain ($30L\times 20L$) with standard reflective boundary conditions to the ones obtained on two different domains of size ($6L\times 4L$) and ($12L\times 8L$) that use the Biot-Savart boundary conditions. We simulate the flow around this heaving airfoil for 100 convective times to ensure that the wake stabilizes, and we gather pressure and viscous forces for the last 20 convective times. Mean lift forces are obtained by time-averaging these forces over the past 20 convective times.
@@
(d) Biot Savart ($12L\times 8L$)
@@
Figure 12: (left) Mean lift force acting on the airfoil at $\text{Re}=100$, for a range of Strouhal numbers for the Biot-Savart and the reflective boundary condition and (right) Poincaré map of the lift and drag coefficients for $St=0.6$
[…]
We expect that in any case where wake history effects are highly dominant, the new Biot-Savart boundary condition will require a domain large enough for this wake to develop “stable” dynamics before exiting the domain. However, we should note that the case of 2D deflected thrust wakes is somewhat pathological, depending on long-range vortex interactions only possible in 2D flows. Indeed, experiments on finite foils have shown that these high-lift deflected jets cannot form for finite wings, see [30, 31].
@@ 6 Conclusion
In this manuscript, we presented a novel boundary condition formulation for external incompressible flow simulations. The method relies on a Biot-Savart integral of the vorticity inside the domain to set the velocity boundary conditions on the domain boundaries. We couple these boundary conditions with an explicit projection scheme to solve the momentum and continuity equation in their primitive variables form. Leveraging the fast decaying properties of the Biot-Savart kernel, we use a multilevel algorithm to reduce the number of operations required to update the boundary nodes from $O(N)$ to $O(\log N)$ and show that the error introduced by this multilevel approach is bounded and can be estimated. We demonstrate that when a body is immersed in a domain using these boundary conditions, a coupling between the immersed body and the domain’s boundary is introduced. This coupling only results in a weakly non-linear pressure Poisson equation and a simple differed correction approach can be used to solve the coupled projection/Biot-Savart step, allowing the use of standard Poisson solver.
@@
With different examples of flow with confined vorticity, we show that the novel boundary conditions allow for a significant reduction in the size of the computational domain without any loss in accuracy. The method is able to exactly capture added-mass force on an accelerated disk with a computational domain only extending $1/2$ diameter away from the plate.
@@
Surprisingly, for vortex street applications, the method demonstrates its robustness and perfectly captures the drag forces and vortex shedding frequencies generated by a 2D circular cylinder at $\text{Re}=200$, even with minimal domain sizes ($5D\times 3D$). Classical reflection boundary conditions require a domain 48 times bigger to generate time-varying forces indistinguishable from that of the domain with our Biot-Savart boundary conditions. Lastly, we show that the method is limited when strong wake history influences the flow in the near wake. For a high-amplitude, high-Strouhal number heaving airfoil, we show that capturing the deflected wake of the flow requires significantly larger domain, implying that the computational advantage of the new boundary condition is reduced.
@@
The generality and accuracy of the new approach, even when simulating flows with vortex wakes that are truncated by the domain boundaries, opens a huge range of potential applications. In addition to the welcomed speed-up enabled by using frankly tiny computational domains, the continued accuracy of the simulations challenges the notion of the importance of far-field wake dynamics on the near-body flow, providing compelling evidence to the contrary.
Appendix A Accounting for symmetries in the Biot-Savart integral
@@
We briefly describe how symmetries can be accounted for in the multilevel evaluation of the Biot-Savart integral Eq. 8. We focus on a description of the 2D case; the extension to the 3D is trivial. For a flow with symmetries, like the 2D flow around a flat plate, see Figure 13, whose axis of symmetries is represented by the dash-dotted line, the computational domain $\Omega$ only models the upper part of the plate. The domain $\Omega^{\dagger}$ is the image of the computational domain $\Omega$ about the symmetry axis.

### v3

@@ 4.2 Wake behind a heaving foil
We immerse a rigid foil of length $L$ moving with velocity $U$ in a viscous fluid. In addition to its steady forward speed, the foil undergoes a pure heave motion of amplitude $h_{0}$
@@
where $h_{0}/L=0.5$ is the non-dimensional amplitude, and the frequency is $f=\text{St}\,U/h_{0}$. For Strouhal number $\text{St}<0.5$, the airfoil generates thrust, and almost no net lift, and for $\text{St}>0.6$, the wake generated by this flapping foil is strongly asymmetric and generates a significant net lift force. We perform simulations of this system for a range of Strouhal numbers and compare the mean lift force and wake obtained on a large domain ($30L\times 20L$) with standard reflective boundary conditions to the ones obtained on two different domains of size ($6L\times 4L$) and ($12L\times 8L$) that use the Biot-Savart boundary conditions. We simulate the flow around this heaving airfoil for 100 convective times to ensure that the wake stabilizes, and we gather pressure and viscous forces for the last 20 convective times. Mean lift forces are obtained by time-averaging these forces over the last 20 convective times.
@@
(d) Biot-Savart ($12L\times 8L$)
@@
Figure 10: (left) Mean lift force acting on the airfoil at $\text{Re}=100$, for a range of Strouhal numbers for the Biot-Savart and the reflective boundary condition and (right) Poincaré map of the lift and drag coefficients for $\text{St}=0.6$
We observe almost no difference between the domain sizes for $\text{St}<0.5$. The wake is symmetric, and the airfoil produces zero mean lift forces; Fig. 10(a). For $\text{St}\geq 0.5$, we observe wake deflection due to the high frequency and large circulation vortices pairing up into vortex-couples as seen in Fig. 9 for $\text{St}=0.6$. Fig. 10(b) shows a Poincaré map for the lift and drag coefficient of the heaving airfoil at $\text{St}=0.6$, with the asymmetry in the lift resulting in a net mean lift coefficient around $C_{L}=-1$ for the very large domain with reflection BCs. The new Biot-Savart boundary conditions predict wake deflection for $\text{St}\geq 0.5$ correctly, but truncating this sensitive vortex wake causes the space between the vortex couples to reduce, Fig. 9(c,d). Using a domain extending only four chord lengths behind the foil results in a 10-40% over-prediction of the measured forces, while doubling the domain recovers the correct spacing and forces to within 5%.
We expect that in any case where wake history effects are highly dominant, the new Biot-Savart boundary condition will require a domain large enough for this wake to develop stable dynamics before exiting the domain. However, we should note that the case of 2D deflected thrust wakes is somewhat pathological, depending on long-range vortex interactions only possible in 2D flows. Indeed, experiments on finite foils have shown that these high-lift deflected jets cannot form for finite wings, see [35, 36].
@@ 5 Conclusion
In this paper, we present a novel coupling to impose unbounded boundary conditions on viscous incompressible flow simulations. The method relies on a Biot-Savart integral of the vorticity inside the domain to set the velocity boundary conditions on the domain boundaries. We couple these boundary conditions with an explicit projection scheme to solve the momentum and continuity equations in their primitive variable form. Critically, we demonstrate that when we apply the Biot-Savart equation to the domain boundaries, a coupling between every point on immersed solid boundaries and the domain boundary is introduced. This coupling results in a dense pressure equation that we solve using iterationss on the projection and Biot-Savart update steps, allowing a standard Poisson solver to be used. This approach also makes the solution extremely robust, enabling simple and high-speed FMM variants to be used without sacrificing solution accuracy. In this work, we use a multi-level clustering method for the $N$ Eulerian-grid vortex sources, and use this to develop a simple Fast Multi-level Method ($\ell$M) for the velocity reconstruction with $O(N)$ complexity that can be run in parallel on CPU and GPU architectures. We show that the error introduced by this multilevel approach is bounded and that its influence on the resulting flow field is minimal because of the iterative partitioned approach used for the pressure.
@@
With different examples of flow with confined vorticity, we show that the novel boundary conditions allow for a significant reduction in the size of the computational domain without any loss in accuracy. The method exactly captures added-mass force on an accelerated disk with a computational domain only extending $1/2$ diameter away from the plate. This domain is also large enough to perfectly replicate the theoretical and previous numerical studies on an impulsively started circular cylinder. In comparison, a domain more than 100 times larger is required when using reflection boundary conditions.
@@
Surprisingly, the method demonstrates its robustness for flows with developed vortex wakes. The developed flow on a sphere is found to be largely independent of the domain size, even for blockage ratios greater than 40%. Lastly, we show that the method is limited when a strong wake history influences the flow in the near wake. For a high-amplitude, high-Strouhal number ($St\geq 0.5$) heaving airfoil, we show that capturing the deflected wake of the flow requires a significantly larger domain, implying that the computational advantage of the new boundary condition is reduced.
@@
The generality and accuracy of the new approach, even when simulating flows with vortex wakes truncated by the domain boundaries, opens a large range of potential applications. In addition to the welcomed speed-up enabled by using minimal computational domains, the continued accuracy of the simulations challenges the notion of the importance of far-field wake dynamics on the near-body flow, providing compelling evidence to the contrary.
Appendix A Multilevel error analysis
A.1 Analytic error bound
We derive the error induced by this multilevel approach, starting with the 2D-case and extending the result to 3D. The velocity at location $\vec{x}$ induced by the circulation from a single cell on level $i+1$ is given by (see Fig. 2)
After rearrangement, the first term equals Eq. 25. The second term is the error due to the multilevel approximation. Its magnitude is
A.2 2D vortex reconstruction example
We next use an analytic 2D vortex field to demonstrate the error in the reconstruction method. The velocity field generated by the Lamb-Chaplygin dipole is defined by the scalar stream function
We reconstruct the velocity field outside the vortex core using the internal vorticity and our multilevel Biot-Savart method and compare that to the analytic result. We reconstruct the velocity on a domain of size $[2^{8},2^{8}]$. The dependency of the absolute error in the reconstructed velocity field on the half-width $\tilde{S}$ of the subdomains are shown in Fig. 11(a). Setting $\tilde{S}$ to the domain size $\cal D$ eliminates the source clustering error, but the result is not exact since the analytic vorticity distribution is sampled on a finite grid of size $\Delta x=0.05R$. This error depends primarily on the distance from the vortex, Fig. 11(a). Progressively reducing the subdomain size introduces the multilevel error quantified in appendix A, which is only weakly dependent on the distance to the vortex. Fig. 11(b) reports the measured computational speed-up achieved by the multilevel method relative to setting $\tilde{S}=\Omega$, demonstrating up to 300$\times$ acceleration.
A.3 3D vortex reconstruction example
and the stream function
with the corresponding errors and speed-up shown in Figure 12. Note that the analytic vorticity of the hill vortex is discontinuous, increasing the truncation error near the vortex compared to the smooth 2D Lamb dipole.
The multilevel error decreases extremely quickly as $\tilde{S}$ increases for three-dimensional flows. This is due to the scaling of the Biot-Savart kernel itself. Similarly, the speed-up is much faster in 3D because each doubling gathers 8 cells instead of 4, with a maximum observed speed-up of nearly 20,000$\times$ using the multilevel method.
Appendix B Accounting for symmetries in the Biot-Savart integral
@@
Symmetries can be accounted for in the multilevel evaluation of the Biot-Savart integral Eq. 8. We focus on a description of the 2D case; the extension to the 3D is trivial. For a flow with symmetries, like the 2D flow around a flat plate, see Figure 13, whose axis of symmetries is represented by the dash-dotted line, the computational domain $\Omega$ only models the upper part of the plate. The domain $\Omega^{\dagger}$ is the image of the computational domain $\Omega$ about the symmetry axis.